# Table-Critic: A Multi-Agent Framework for Collaborative Criticism and Refinement in Table Reasoning

### v1

@@ Abstract
Despite the remarkable capabilities of large language models (LLMs) in various reasoning tasks, they still struggle with table reasoning tasks, particularly in maintaining consistency throughout multi-step reasoning processes. While existing approaches have explored various decomposition strategies, they often lack effective mechanisms to identify and correct errors in intermediate reasoning steps, leading to cascading error propagation. To address these issues, we propose Table-Critic, a novel multi-agent framework that facilitates collaborative criticism and iterative refinement of the reasoning process until convergence to correct solutions. Our framework consists of four specialized agents: a Judge for error identification, a Critic for comprehensive critiques, a Refiner for process improvement, and a Curator for pattern distillation. To effectively deal with diverse and unpredictable error types, we introduce a self-evolving template tree that systematically accumulates critique knowledge through experience-driven learning and guides future reflections. Extensive experiments have demonstrated that Table-Critic achieves substantial improvements over existing methods, achieving superior accuracy and error correction rates while maintaining computational efficiency and lower solution degradation rate.
@@ 1 Introduction
Despite significant advances in various reasoning tasks (Plaat et al., 2024; Yu et al., 2024; Chen et al., 2024a, b; Guo et al., 2025), large language models (LLMs) (Yang et al., 2024a; Dubey et al., 2024; Anthropic, 2024; Mesnard et al., 2024; Hurst et al., 2024) face substantial challenges in handling semi-structured data, such as table reasoning tasks, as they require both understanding of tabular structures and precise localization of relevant entries in redundant and noisy information (Zhao et al., 2024; Chen et al., 2024c; Zhang et al., 2025).
@@ 3.1 Overview
To effectively implement the human-like correction process in multi-step reasoning, we propose a collaborative multi-agent framework, Table-Critic. As illustrated in Figure 1, this framework decomposes the complex reasoning refinement task into four specialized functions: error detection (Judge), critique generation (Critic), reasoning refinement (Refiner), and experience learning (Curator). These agents work in concert to progressively improve reasoning quality while accumulating valuable correction experiences. Specifically, given a table $\mathbb{T}$ and a question $q$, these agents iteratively refine the initial reasoning chain $\tau=\{s_{1},s_{2},...,s_{n}\}$ until reaching a satisfactory solution.§§§We use Chain-of-Table (Wang et al., 2024) for initial chains, though our framework is applicable to other methods. The refinement process is guided by a self-evolving template tree $\mathcal{T}$ that systematically accumulates critique patterns from past experiences.
@@ Appendix E Prompts and Case Study
This appendix provides comprehensive instructions and illustrative examples for three intelligent agents: the Judge Agent, the Critic Agent, and the Refiner Agent. These agents are designed to collaboratively evaluate and refine reasoning processes applied to table-based questions. Figures 5 and 6 offer detailed guidance for the Judge Agent, including step-by-step procedures to assess the validity of reasoning steps, pinpoint errors, and categorize conclusions (e.g., correct, incorrect with identified error route, or random error). Figures 7 and 8 explain how the Critic Agent systematically evaluates each reasoning step, highlights the first incorrect step, and provides constructive critiques. Additionally, Figure 9 introduces the Refiner Agent, demonstrating how critiques are utilized to refine reasoning steps, ensuring accurate and complete solutions.

### v2

@@ Abstract
Despite the remarkable capabilities of large language models (LLMs) in various reasoning tasks, they still struggle with table reasoning tasks, particularly in maintaining consistency throughout multi-step reasoning processes. While existing approaches have explored various decomposition strategies, they often lack effective mechanisms to identify and correct errors in intermediate reasoning steps, leading to cascading error propagation. To address these issues, we propose Table-Critic, a novel multi-agent framework that facilitates collaborative criticism and iterative refinement of the reasoning process until convergence to correct solutions. Our framework consists of four specialized agents: a Judge for error identification, a Critic for comprehensive critiques, a Refiner for process improvement, and a Curator for pattern distillation. To effectively deal with diverse and unpredictable error types, we introduce a self-evolving template tree that systematically accumulates critique knowledge through experience-driven learning and guides future reflections. Extensive experiments have demonstrated that Table-Critic achieves substantial improvements over existing methods, achieving superior accuracy and error correction rates while maintaining computational efficiency and lower solution degradation rate. The code is available at https://github.com/Peiying-Yu/Table-Critic.
@@ 1 Introduction
Despite significant advances in various reasoning tasks (Plaat et al., 2024; Yu et al., 2024; Chen et al., 2024a, b; Guo et al., 2025; Xu et al., 2025), large language models (LLMs) (Yang et al., 2024a; Dubey et al., 2024; Anthropic, 2024; Mesnard et al., 2024; Hurst et al., 2024) face substantial challenges in handling semi-structured data, such as table reasoning tasks, as they require both understanding of tabular structures and precise localization of relevant entries in redundant and noisy information (Zhao et al., 2024; Chen et al., 2024c; Zhang et al., 2025).
@@ 3.1 Overview
To effectively implement the human-like correction process in multi-step reasoning, we propose a collaborative multi-agent framework, Table-Critic. As illustrated in Figure 1, this framework decomposes the complex reasoning refinement task into four specialized functions: error detection (Judge), critique generation (Critic), reasoning refinement (Refiner), and experience learning (Curator). These agents work in concert to progressively improve reasoning quality while accumulating valuable correction experiences. Specifically, given a table $\mathbb{T}$ and a question $q$, these agents iteratively refine the initial reasoning chain $\tau=\{s_{1},s_{2},...,s_{n}\}$ until reaching a satisfactory solution. [^1] We use Chain-of-Table (Wang et al., 2024) for initial chains, though our framework is applicable to other methods. The refinement process is guided by a self-evolving template tree $\mathcal{T}$ that systematically accumulates critique patterns from past experiences.
@@ 5 Conclusion
Acknowledgments
This work was supported by three NSFC grants, i.e., No.62006166, No.62376178 and No.62076175. This work was also supported by Collaborative Innovation Center of Novel Software Technology and Industrialization, and supported by a Project Funded by the Priority Academic Program Development of Jiangsu Higher Education Institutions (PAPD).
@@ Appendix E Prompts and Case Study
This appendix provides comprehensive instructions and illustrative examples for four intelligent agents: the Judge Agent, the Critic Agent, the Refiner Agent, and the Curator Agent. These agents are designed to collaboratively evaluate and refine reasoning processes applied to table-based questions.
Figures 5 and 6 offer detailed guidance for the Judge Agent, including step-by-step procedures to assess the validity of reasoning steps, pinpoint errors, and categorize conclusions (e.g., correct, incorrect with identified error route, or random error). Figures 7 and 8 explain how the Critic Agent systematically evaluates each reasoning step, highlights the first incorrect step, and provides constructive critiques. Figure 9 introduces the Refiner Agent, demonstrating how critiques are utilized to refine reasoning steps, ensuring accurate and complete solutions.
As for Curator Agent, Curator expands the template tree $\mathcal{T}$ by adding templates to existing leaf nodes or expanding tree branches (as described in Algorithm 2). This process begins by invoking the Judge agent to determine the routing path and evaluation status of the refinement history within the context of the current template tree (Line 1). The outcome, denoted as $route$, indicates whether the refinement aligns successfully with any existing route in $\mathcal{T}$.
When the $route$ is successful within $\mathcal{T}$ ($route$.status = SUCCESS), Curator will perform Adding Templates or Vertical Expansion. To make this decision, Curator uses a prompt to assess the similarity between the refinement history $H$ and the sampled templates $\mathcal{T}_{r}$ (see Figure 10). The agent then checks whether the refinement history $H$ is sufficiently similar to the node $T_{r}$ (Line 4). If so (Figure 11), the template tree is directly augmented by adding the new template derived from $H$ through the AddTemplate function (Line 5). If the refinement history diverges in content but follows a related path (Figure 12), a VerticalExpansion is performed to hierarchically extend the tree structure while preserving the routing path (Line 7).
Conversely, if the refinement path does not correspond to any existing route ($route$.status $\neq$ SUCCESS), the agent invokes HorizontalExpansion to create a new sibling branch in the template tree, thereby broadening its representational capacity (Line 10). This horizontal expansion process is shown in Figure 13, where a new branch is added to accommodate unmatched templates.
This process ensures that the template tree $\mathcal{T}$ evolves into a more expressive and comprehensive structure $\mathcal{T}^{\prime}$, capable of supporting a wider range of reasoning refinements. The final evolved tree $\mathcal{T}^{\prime}$ is returned as the output (Line 12), encapsulating the incremental knowledge acquired through iterative refinement.
Algorithm 2 The implementation of Curator Agent
0:  The template tree $\mathcal{T}$, the refinement history $H$.
0:  The evolved tree $\mathcal{T}^{\prime}$.
1:  $route\leftarrow\texttt{Judge}(\mathcal{T},H)$
2:  if $route.\text{status}=\text{SUCCESS}$ then
3:     $\mathcal{T}_{r}\leftarrow$ Sample Templates using $route$ in the $\mathcal{T}$
4:     if $H$ is similar to $\mathcal{T}_{r}$ then
5:        $\mathcal{T}^{\prime}\leftarrow\texttt{AddTemplate}(\mathcal{T},H)$
6:     else
7:        $\mathcal{T}^{\prime}\leftarrow\texttt{VerticalExpansion}(\mathcal{T},H)$
8:     end if
9:  else
10:     $\mathcal{T}^{\prime}\leftarrow\texttt{HorizontalExpansion}(\mathcal{T},H)$
11:  end if
12:  return  Evolved template tree $\mathcal{T}^{\prime}$
@@
You are organizing hierarchical categories and their associated few-shot examples. Currently, you have two lists of few-shot examples under the same category. Your task is to decide whether these two lists can be meaningfully split into two distinct subcategories. Instructions: Analyze the examples in the two lists to determine if there is a clear and meaningful distinction between them. - If a distinction exists, create two subcategories and assign each list to one of them. - If no clear distinction exists, retain both lists under the original parent category. Provide a clear explanation for your decision on whether to split or merge the lists, based on their content. Parent Category: row error List 1: [ (Template 1), (Template 2), ... ] List 2: [ (New Template) ]
Figure 10: Instructions for determining whether the refinement history $H$ is similar to sampled templates $\mathcal{T}_{r}$. It guides the model to decide if the new templates should be merged with the existing category or split into distinct subcategories.
Explanation: ... Determination: List 1: <row error> List 2: <row error>
Figure 11: Model explanation and categorization result when determining whether to add the refinement history $H$ to existing templates $\mathcal{T}_{r}$. This corresponds to the AddTemplate operation in Algorithm 2.
Explanation: ... Determination: List 1: <row misidentification error> List 2: <row omission error>
Figure 12: Model explanation and categorization decision when refinement history $H$ is not exactly similar to existing templates $\mathcal{T}_{r}$. This corresponds to the VerticalExpansion step in Algorithm 2.
You are given an error tree represented as a dictionary and a template describing a specific error. If the error path corresponding to the template cannot be found in the error tree, extend the error tree by adding a new branch at the appropriate location. Ensure that: - The new branch aligns logically with the existing structure of the error tree. - The template can be correctly integrated into the tree under the newly added branch. Error Tree: { "sub-table error": { "row error": "<END>", "column error": "<END>" } } Template: (New Template) ------------------------------------------------------------------------------------------ For example, if the return result is as follows: Addition: (final query error -> <END>) Curator will add a "final query error" node under the root node for this new template.
Figure 13: Instructions and example of HorizontalExpansion. Prompt guiding the model to extend the template tree horizontally by adding a new branch when the refinement history $H$ does not match any existing template paths. This corresponds to the HorizontalExpansion step in Algorithm 2.